# Enhancing LLMs for Power System Simulations: A Feedback-driven Multi-agent Framework

### v1

@@ IV Environmental Acting Module with Feedback
Despite the reinforcement brought by the enhanced RAG and reasoning modules, the coding agent may still encounter errors during simulation code generation. To address this, it is essential to enable direct interaction between the LLM and the simulation environment, allowing the agent to receive execution feedback and iteratively refine its code. To this end, we propose an environmental acting module with an error feedback mechanism that integrates with both the RAG and reasoning modules, as illustrated in Fig. 5. The components of this module are described in the following.
@@ IV-A Code Execution and Detection
The simulation code, generated by the coding agent from the enhanced reasoning module, is executed using the simulation environment API connected to a specific power systems simulation tool. Following execution, the simulation environment produces results, which are then checked for error signals. Specifically:
@@ IV-B Error Handling and Feedback Loop
Figure 5: Environmental acting module with an error feedback mechanism that integrates with both the RAG and reasoning modules.
@@
(GPT4o: API of gpt-4o-2024-05-13; CGPT4o: ChatGPT4o Web Interface; o1p: O1-preview Web Interface)
GPT4o Full GPT4o PR GPT4o RSR GPT4o SR GPT4o Sole GPT4o NC GPT4o NP GPT4o NS GPT4o NR GPT4o NCS GPT4o RSRNW CGPT4o R o1p Sole
@@ IV-C Enhanced RAG and Reasoning Module Interplay
The error report and feedback are then processed as a new request by the retrieval agent in the enhanced RAG module. This agent retrieves relevant information based on the error report (the query planning can also be applied to error reporting by simply replacing the identification of function/option keywords with the identification of error-related keywords). The retrieved information is then passed to the enhanced reasoning module, where the coding agent incorporates both the retrieval results and the correction request to revise the simulation code. The loop proceeds until the code meets both requirements, or until the stopping criterion is satisfied.
@@ V Case Study
To comprehensively validate the proposed framework, we carry out a range of tests differing in three key dimensions: (i) distinct combinations of the proposed strategies within the framework to evaluate each strategy’s independent effectiveness; (ii) different simulation environments, specifically Daline [23] and MATPOWER [24], which include tools both familiar and unfamiliar to LLMs [^1] Daline is available [here] with a user manual in [25]. MATPOWER is available [here] with a user manual in [26]., to demonstrate the framework’s versatility across various applications; and (iii) a wide array of simulation tasks, spanning normal to complex scenarios, to assess the framework’s performance across various simulation demands.
The following sections detail the case study configurations, followed by an analysis of the simulation outcomes for Daline and MATPOWER. Eventually, the cost of using LLMs to perform power system simulations is discussed.
@@ V-A Settings
Firstly, Table I presents the distinct combinations of the proposed strategies within the framework employed in the evaluation, with the proposed strategies shaded in gray.
@@
Thirdly, 34 simulation tasks have been used to test the framework with Daline, comprising 7 complex tasks and 27 standard tasks. Similarly, for MATPOWER, 35 simulation tasks have been defined, including 8 complex tasks and 27 standard tasks. These tasks comprehensively cover the functionalities of both simulation tools, aiming to include the majority of available options. A selection of representative tasks is given in Fig. 6, intended as an exemplary overview.
@@
Finally, for the performance evaluation, each scheme is allowed up to $N_{max}$ attempts of addressing a simulation task (with $N_{max}=3$ for Daline and $N_{max}=5$ for MATPOWER, as the latter is relatively more complex). A scheme earns 100 points per attempt for exactly correct simulation results without irrelevant settings in the code, 50 points for correct simulation results but with irrelevant settings in the code, and 0 points for any incorrect simulation outcomes. Subsequent attempts are only made if the previous attempt encounters execution errors, and any unused attempts receive the same score as the last attempt. The success rate for each scheme is defined as the total points earned divided by the maximum possible score, resulting in a success rate between 0% and 100%.
@@ V-B Evaluation on Daline
Figure 7: Distribution of scores achieved across attempts for each evaluated scheme, separated by complex and standard tasks (Simulation Environment: Daline).
Figure 8: Success rates for each scheme, itemized by all tasks combined, complex tasks only, standard tasks only, as well as for the first attempt success rate and the final attempt success rate (Simulation Environment: Daline).
@@
While equipped with environmental interaction and feedback mechanisms, GPT4o-Sole still demonstrates a 0% success rate for both complex and standard tasks, indicating that GPT4o has not previously encountered Daline. Even with a complete knowledge base supported by RAG — either through the standard RAG or OpenAI’s official RAG — the resulting schemes, GPT4o-SR and CGPT4o-R, achieve success rates of only 31.37% and 33.82% across all tasks, respectively. This suggests that, even with RAG support, the latest language model, GPT4o, still lacks reliable performance in simulations. In contrast, the scheme equipped with our proposed full framework, GPT4o-Full, achieves a success rate of 93.13% across all tasks — a significant improvement that highlights the effectiveness of the proposed framework.
@@
The high success rate of 93.13% over all taks for GPT4o-Full is due to the cumulative effects of using the complete proposed framework. Comparing other schemes with GPT4o-Full gives an indication of the impact of omitted strategies. For instance, although GPT4o-NP includes most reasoning enhancement strategies, it lacks the triple-based structured option document, resulting in a reduced success rate of 81.37%. On the other hand, omitting the few-shot CoT for reasoning, as in GPT4o-NCS, lowers success to 65.19%. When few-shot CoT is employed, but the proposed query planning is omitted, as in GPT4o-RSR, the success rate for complex tasks drops to 66.67%, particularly due to the deteriorated performance for the complex tasks. While similar comparisons can be drawn across all schemes, our goal here is not to argue which strategy provides the highest improvement, but to emphasize that high success relies on the combined effect of multiple strategies.
@@
In general, more complex tasks — those with multiple sub-requests — tend to increase the likelihood of errors in LLMs, resulting in lower success rates compared to standard tasks, as observed across most schemes. However, with our proposed full framework, GPT4o-Full, the performance gap between complex and standard tasks narrows significantly, as shown in both the score distributions in Fig. 7 and the success rates in Fig. 8. This suggests that, with enhanced reasoning capabilities and the more effective RAG design, GPT4o-Full effectively identifies and addresses the sub-requests within complex tasks, similar to how it handles standard tasks. This enables LLMs to better manage complex tasks.
@@
The comparison between the first-attempt and final-attempt success rates demonstrates the effectiveness of environmental interaction and feedback mechanisms. As shown in Fig. 8, the final-attempt success rate is always higher than the first-attempt rate, particularly for schemes that incorporate fewer strategies from our framework. These schemes typically have either reduced reasoning capability or limited retrieval information, making environmental interaction and feedback crucial for error correction. However, for GPT4o-Full, the difference between first-attempt and final-attempt success rates is relatively small, as GPT4o-Full often completes the Daline simulation task successfully on the first attempt. This further highlights the effectiveness of the proposed framework. One noteworthy point is that the effectiveness of automatic error correction is partly influenced by the quality of the simulation tool’s error-reporting system — whether it provides clear, code-specific error messages. This feature affects the LLM’s capability to interpret and resolve issues in the generated code. In the absence of such a feature, as with GPT4o-RSRNW, the success rate drops to 78.43%, with negligible improvement between the first and final attempts. This underscores that without a well-developed error-reporting system, iterative refinement may yield limited benefit. Although tools like Daline and MATPOWER include well-developed error reporting (thereby achieving high correction accuracy), for other simulation tools where such systems are less developed, reinforcing them is recommended. In fact, our framework also enables LLMs to detect vulnerabilities within a simulation tool’s error-reporting system, particularly when mistakes occur. These errors may not stem from limitations of the LLMs or the framework itself but rather from inherent design issues within the tools.
@@
Figure 9: Scores achieved by each evaluated scheme in individual attempts when handling complex tasks. The automatic error correction mechanism aids schemes equipped with our framework, such as GPT4o-PR, in correcting their behavior when encountering errors in initial attempts.
@@
Figure 10: Distribution of scores achieved across attempts for each evaluated scheme, separated by complex and standard tasks (Simulation Environment: MATPOWER).
@@
Figure 11: Success rates for each scheme, broken down by all tasks combined, complex tasks only, standard tasks only, as well as for the first attempt success rate and the final attempt success rate (Simulation Environment: MATPOWER; o1p-Sole is only tested by complex tasks).

### v2

@@ IV Environmental Acting Module with Feedback
Despite the reinforcement brought by the enhanced RAG and reasoning modules, the coding agent may still encounter errors during simulation code generation. To address this, it is essential to enable direct interaction between the LLM and the simulation environment. It allows the agent to receive execution feedback and iteratively refine its code. To this end, this paper proposes an environmental acting module with an error-feedback mechanism that integrates with both the RAG and reasoning modules, as illustrated in Fig. 5. The components of this module are described in the following.
@@ IV-A Code Execution and Detection
The simulation code, generated by the coding agent from the enhanced reasoning module, is executed using the simulation environment API connected to a specific power system simulations tool. Following execution, the simulation environment produces results, which are then checked for error signals. Specifically:
@@ IV-B Error Handling and Feedback Loop
Figure 5: Environmental acting module with an error-feedback mechanism that integrates with both the RAG and reasoning modules.
@@
(GPT4o: APIs of gpt-4o-2024-05-13 and gpt-4o-2024-08-06, with the exact version specified in each evaluation; CGPT4o: ChatGPT4o; o1p: o1-preview)
GPT4o Full GPT4o PR GPT4o RSR GPT4o SR GPT4o Sole GPT4o NC GPT4o NP GPT4o NS GPT4o NR GPT4o NCS GPT4o RSRNW CGPT4o R o1p Sole GPT4o Sole-SFT
@@ IV-C Enhanced RAG and Reasoning Module Interplay
The error report and feedback are then processed as a new request by the retrieval agent in the enhanced RAG module. This agent retrieves relevant information based on the error report (the query planning can also be applied to error reporting by simply replacing the identification of function/option keywords with the identification of error-related keywords). The retrieved information is then passed to the enhanced reasoning module. The coding agent there uses both the retrieval results and the correction request to revise the simulation code. The updated code is fed back into the environmental acting module. This loop continues until the code meets all requirements or the stopping criterion is reached.
This design achieves two main objectives. First, the system continuously monitors simulation executions and triggers adaptive adjustments as needed, ensuring that all module interdependencies are effectively managed. Second, it handles failure modes. The error-feedback mechanism detects failures in both the RAG and reasoning modules—failures that can lead to simulation errors. Once a failure is detected, the mechanism initiates iterative corrective actions using adaptive prompts, which trigger updated retrievals and adjustments, thereby dynamically managing module failures. This conclusion is further reinforced by the case studies in Section V.
IV-D Adaptability
The environmental acting module features a tool-independent design. First, it leverages the error detection and reporting systems common to many power system simulation tools. Second, it formulates the error-feedback loop in a general manner, as detailed in Section IV-B. Each component is independent of any specific simulation tool. This design not only supports automatic error correction but also improve adaptability to diverse simulation platforms.
IV-E Pseudocode for the Full Framework
Algorithm 2 presents the complete pseudocode for the proposed framework. The algorithm employs self-explanatory pseudo-functions and pseudo-options to illustrate the core modules, including their internal processes and the interplay among them.
Input:  modelVer: Model version from Table I
config: Configuration settings from Table II
ragPrompt: Query planning prompt from Fig. 2
reasonPrompt: Structured reasoning prompt from Fig. 4
errTemplate: Error report template from Fig. 5
vectorDB: Vector database from Algorithm 1
task: Simulation task
maxAttempts: Max attempt number
apiKey: Credentials for LLMs
returnNum: Number of top-relevant chunks to return
Output:  simResult: Simulation result
code: Generated code
1
// Preparation;
2 attemptTime $\leftarrow$ 0;
3 chatHistory $\leftarrow$ emptySet;
4 env $\leftarrow$ activateSimulationEnvironment()
// Enhanced RAG Module;
5 retrievalLLM $\leftarrow$ setupLLM(config, modelVer, ragPrompt, apiKey);
6 keywords $\leftarrow$ generateKeywords(retrievalLLM, task);
7 retrievalInfo $\leftarrow$ parallelRetrieval(retrievalLLM, keywords, vectorDB, returnNum);
8
// Enhanced Reasoning Module;
9 reasonPrompt $\leftarrow$ insertRetrieval(reasonPrompt, retrievalInfo);
10 codeLLM $\leftarrow$ setupLLM(config, modelVer, reasonPrompt, apiKey);
11 [code, chatHistory] $\leftarrow$ generateCode(codeLLM, task, chatHistory);
12
// Environmental Acting Module with Feedback;
13 [simResult, err] $\leftarrow$ runSimulation(code, env);
14
15while err $\neq$ null and attemptTime $\leq$ maxAttempts do
16       errReport $\leftarrow$ generateErrorReport(code, err, errTemplate);
17       keywords $\leftarrow$ generateKeywords(retrievalLLM, errReport);
18       newRetrievalInfo $\leftarrow$ parallelRetrieval(retrievalLLM, keywords, vectorDB, returnNum);
19       reasonPrompt $\leftarrow$ insertRetrieval(reasonPrompt, newRetrievalInfo);
20       codeLLM $\leftarrow$ setupLLM(config, modelVer, reasonPrompt, apiKey);
21       [code, chatHistory] $\leftarrow$ generateCode(codeLLM, errReport, chatHistory);
22       [simResult, err] $\leftarrow$ runSimulation(code, env);
23       attemptTime $\leftarrow$ attemptTime + 1;
24
25
26if err $\neq$ null then
27       reportFailure(code);
28       saveFailedCode(code);
29
30else
31       outputResult(simResult);
32       saveSimulationCode(code);
33
Algorithm 2 Feedback-driven Multi-agent Framework
Remark: It is important to emphasize that the proposed framework is built on the premise that simulation tools are learnable by humans through sufficient documentation. If a human can use a simulation tool based on its manuals, then the proposed integrated RAG, reasoning, and feedback modules can replicate and enhance this process to improve efficiency. Hence, for scenarios where documentation is absent, it is suggested to generate necessary documents, possibly with LLMs’ help. These documents can then be incorporated into the proposed framework.
@@ V Case Study
To comprehensively validate the proposed framework, a range of tests have been carried out, differing in three key dimensions: (i) distinct combinations of the proposed strategies within the framework to evaluate each strategy’s independent effectiveness; (ii) different simulation environments, specifically Daline [23] and MATPOWER [24], which include tools both familiar and unfamiliar to LLMs [^1] Daline is available [here] with a user manual in [25]. MATPOWER is available [here] with a user manual in [26]., to demonstrate the framework’s versatility across various applications; and (iii) a wide array of simulation tasks, spanning normal to complex scenarios, to assess the framework’s performance across various simulation demands.
The following sections detail the case study configurations, followed by an analysis of the simulation outcomes for Daline and MATPOWER. Eventually, the cost of using LLMs to perform power system simulations is discussed. All supporting materials, including prompts, knowledge bases, simulation tasks, training datasets, and results (i.e., the generated codes and their benchmarks, totaling 870 coding files) will be available upon the acceptance of this paper.
@@ V-A Settings
Firstly, Table I presents the distinct combinations of the proposed strategies within the framework employed in the evaluation, with the proposed strategies shaded in gray. The model versions of LLMs can also be found in the caption of Table I, as well as in the caption of each evaluation figure/table. Meanwhile, the configurations used in all the evaluations are detailed in Table II.
TABLE II: Configuration settings for evaluations (for the LLM model versions, please see Table I or the captain of each evaluation figure/table)
RAG Configuration
Parameter Value Description
Chunk size for .txt 30 Number of words per chunk for text files
Chunk size for .pdf 50 Number of words per chunk for PDF files
Embedding Model v1 See [here] for the model text-embedding-v1
Return number 20 Number of top-relevant chunks to return
LLM Configuration
Parameter Value Description
Temperature 0.1 Randomness (lower = more predictable)
Max Tokens 4096 Max. length of the output in tokens
Top P 1 Probability threshold for token selection
Frequency Penalty 0 Zero means common words are not suppressed
Presence Penalty 0 Zero allows natural repetition when needed
Acting Configuration
Parameter Value Description
Daline $N_{\max}^{t}$ 3 Max. attempt number for Daline task $t$
MATPOWER $N_{\max}^{t}$ 5 Max. attempt number for MATPOWER task $t$
@@
Thirdly, 34 simulation tasks were used to evaluate the proposed framework on Daline. Similarly, for MATPOWER, 35 simulation tasks have been defined, including 8 complex tasks and 27 standard tasks.
  • The simulation tasks of Daline span from basic data generation to advanced workflows that integrate data creation, data corruption, noise handling, outlier filtering, model training, method comparison, result visualization, and full-cycle management. The simulation suite exercises all distinct functions in Daline across various power system cases. The tasks employ multiple modeling methods (e.g., Least Squares with Huber Weighting Function, Partial Least Squares with Clustering, Ridge Regression with Voltage-angle Coupling, Locally Weighted Ridge Regression, State-independent Voltage-angle Decoupled Method with Data Correction, to name a few) and a wide array of parameters (sample sizes, base types, noise levels, outlier techniques, cross-validation settings, method hyperparameters, etc.).
  • The simulation tasks of MATPOWER span standard AC and DC optimal power flow (or just load flow) analyses as well as advanced continuation power flow (CPF) studies. They employ a variety of solvers and algorithms—including Newton-Raphson, Fast-Decoupled (both XB and BX versions), Gauss-Seidel, Implicit Z-bus Gauss, MIPS, fmincon, and GUROBI—across numerous test cases such as the IEEE 9-, 14-, 30-, 57-, 118-, and 145-bus systems, along with specialized cases like the 4-bus, 6-bus, 39-bus New England, 51-bus radial, 60-bus Nordic, and IEEE RTS 24-bus systems. Key parameters such as maximum iterations, mismatch tolerances, coordinate representations, and branch flow constraints were systematically varied. In the CPF analyses, both natural and pseudo arc length parameterizations were explored with adaptive step sizing, explicit nose point detection, and generation/load scaling. Additional configurations addressed generator reactive power limits, voltage setpoints, alternative formulations (current versus power balance), and solver-specific settings like gradient, optimality, and termination tolerances.
Overall, these tasks comprehensively cover both basic and advanced functionalities of Daline and MATPOWER across a broad range of power system analysis scenarios. A selection of representative tasks is given in Fig. 6, intended as an exemplary overview.
@@
Finally, to evaluate each scheme’s performance, we compute a success rate based on the points obtained across multiple simulation tasks. Let $T$ denote the total number of simulation tasks. For each task $t$ ($1\leq t\leq T$), let $N_{\max}^{(t)}$ be the maximum number of allowed attempts for that task and let $P_{t,i}$ be the points awarded on attempt $i$ (with $1\leq i\leq N_{\max}^{(t)}$). The scoring for task $t$ at attempt $i$ is defined as follows:
  • $P_{t,i}=100$ points if the simulation result is exactly correct with no irrelevant settings in the code.
  • $P_{t,i}=50$ points if the simulation result is correct but contains irrelevant settings.
  • $P_{t,i}=0$ points if the simulation result is incorrect.
Subsequent attempts are made only if the previous attempt resulted in an error, and any unused attempts are assigned the same score as the last executed attempt. Thus, the total score for task $t$ is given by:
$S_{t}=\sum_{i=1}^{N_{\max}^{(t)}}P_{t,i},$
with a maximum possible score of $100\times N_{\max}^{(t)}$ for that task. Finally, the overall success rate across all tasks is computed as:
$R_{\text{overall}}=\frac{\sum_{t=1}^{T}S_{t}}{\sum_{t=1}^{T}\left(100\times N_{\max}^{(t)}\right)}\times 100\%.$
In the following experiments, $N_{\max}^{t}=3$ ($\forall t$) is set for Daline tasks, and $N_{\max}^{t}=5$ ($\forall t$) is set for MATPOWER tasks, reflecting the increased complexity of the latter.
@@ V-B Evaluation on Daline
While equipped with environmental interaction and feedback mechanisms, GPT4o-Sole still demonstrates a 0% success rate for both complex and standard tasks, indicating that GPT4o has not previously encountered Daline. Even with a complete knowledge base supported by RAG — either through the standard RAG or OpenAI’s official RAG — the resulting schemes, GPT4o-SR and CGPT4o-R, achieve success rates of only 31.37% and 33.82% across all tasks, respectively. This suggests that, even with RAG support, the latest language model, GPT4o, still lacks reliable performance in simulations. In contrast, the scheme equipped with the proposed full framework, GPT4o-Full, achieves a success rate of 93.13% across all tasks — a significant improvement that highlights the effectiveness of the proposed framework.
@@
The high success rate of 93.13% over all taks for GPT4o-Full is due to the cumulative effects of using the complete proposed framework. Comparing other schemes with GPT4o-Full gives an indication of the impact of omitted strategies. For instance, although GPT4o-NP includes most reasoning enhancement strategies, it lacks the triple-based structured option document, resulting in a reduced success rate of 81.37%. On the other hand, omitting the few-shot CoT for reasoning, as in GPT4o-NCS, lowers success to 65.19%. When few-shot CoT is employed, but the proposed query planning is omitted, as in GPT4o-RSR, the success rate for complex tasks drops to 66.67%, particularly due to the deteriorated performance for the complex tasks. Similar comparisons can be drawn across all schemes. However, the goal is not to determine which strategy provides the highest improvement. Instead, this paper aims to emphasize that high success relies on the combined effect of multiple strategies.
@@
In general, more complex tasks — those with multiple sub-requests — tend to increase the likelihood of errors in LLMs, resulting in lower success rates compared to standard tasks, as observed across most schemes. However, with the proposed full framework, GPT4o-Full, the performance gap between complex and standard tasks narrows significantly, as shown in both Fig. 7 and Fig. 8. This suggests that, with enhanced reasoning capabilities and the more effective RAG design, GPT4o-Full effectively identifies and addresses the sub-requests within complex tasks, similar to how it handles standard tasks. This enables LLMs to better manage complex tasks.
Figure 7: Distribution of scores achieved across attempts for each evaluated scheme (simulation environment: Daline; GPT4o version: gpt-4o-2024-05-13).
Figure 8: Success rates for each scheme, itemized by all tasks combined, complex tasks only, standard tasks only, as well as for the first attempt success rate and the final attempt success rate (simulation environment: Daline; GPT4o version: gpt-4o-2024-05-13).
@@
The comparison between the first-attempt and final-attempt success rates demonstrates the effectiveness of environmental interaction and feedback mechanisms. As shown in Fig. 8, the final-attempt success rate is always higher than the first-attempt rate, particularly for schemes that incorporate fewer strategies from the proposed framework. These schemes typically have either reduced reasoning capability or limited retrieval information, making environmental interaction and feedback crucial for error correction. However, for GPT4o-Full, the difference between first-attempt and final-attempt success rates is relatively small, as GPT4o-Full often completes the Daline simulation task successfully on the first attempt. This further highlights the effectiveness of the proposed framework. One noteworthy point is that the effectiveness of automatic error correction depends partly on the quality of the simulation tool’s error-reporting system. Specifically, it matters whether the system provides clear, code-specific error messages. This feature affects the LLM’s capability to interpret and resolve issues in the generated code. In the absence of such a feature, as with GPT4o-RSRNW, the success rate drops to 78.43%, with negligible improvement between the first and final attempts. This underscores that without a well-developed error-reporting system, iterative refinement may yield limited benefit. Although tools like Daline and MATPOWER include well-developed error reporting (thereby achieving high correction accuracy), for other simulation tools where such systems are less developed, reinforcing them is recommended. In fact, the proposed framework also enables LLMs to detect vulnerabilities within a simulation tool’s error-reporting system, particularly when mistakes occur. These errors may not stem from limitations of the LLMs or the framework itself but rather from inherent design issues within the tools.
Enhancing error-reporting systems in simulation tools is beyond this paper’s scope. However, for tools with ambiguous error messages, the following is proposed:
  • If internal modifications are feasible, a layered checking mechanism can be implemented, where an agent evaluates error reports across functional layers, identifies vulnerabilities, and iteratively refines messages.
  • If internal modifications are not feasible, three external strategies may be employed: (i) For LLMs with long-context capabilities, sending the entire underlying code for analysis is possible but costly. (ii) A more efficient alternative is to batch extract relevant code segments—from surface-layer to deeper-layer functions—and submit them alongside error messages to the agents for analysis. (iii) An active trial-and-error mechanism based on agents can automatically test code with intentional bugs, mapping observed errors to ambiguous messages, in order to flag problematic outputs and facilitate learning.
These strategies can be seamlessly integrated into the proposed framework. Internal modifications require no changes to the framework since they occur within the simulation tool. Notably, the layered checking mechanism has already been implemented in Daline, proving highly effective not only for the proposed framework but also as a robust enhancement for Daline itself. For external strategies, the only adjustment needed is to extend the error report with relevant code extraction or error-message mapping. Future research will focus on leveraging the LLM-based trial-and-error mechanism to systematically map ambiguous error messages, reducing diagnosis overhead and improving robustness.
@@
Figure 9: Scores achieved by each evaluated scheme in individual attempts when handling complex tasks. The automatic error correction mechanism aids schemes in correcting their behavior when encountering errors in initial attempts (simulation environment: MATPOWER; GPT4o version: gpt-4o-2024-05-13).
@@
Figure 10: Distribution of scores achieved across attempts for each evaluated scheme, separated by complex and standard tasks (simulation environment: MATPOWER; GPT4o version: gpt-4o-2024-05-13).
@@
Figure 11: Success rates for each scheme, broken down by all tasks combined, complex tasks only, standard tasks only, as well as for the first attempt success rate and the final attempt success rate (simulation environment: MATPOWER; GPT4o version: gpt-4o-2024-05-13; o1p-Sole is only tested by complex tasks).